Systems Staff Meeting Notes 4/24/00

Systems Staff Meeting Notes
April 24, 2000

Present: Will Chaney, Bill Covey, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Martha Hruska presented Service Award pins and certificates to the following Systems personnel:
    Bobbie Meng - 25 years
    Bill Covey - 15 years
    Winston Harris - 5 years
    Everett Phillips - 5 years
    Congratulations!
  2. Will went to Busch Gardens over this past weekend & had such a good time he's returning this weekend!
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders will go out for the switches and wiring to replace the hubs. The point where the switches hook up to the network will have to be moved. We'll have approximately 70 total pulls. The advantage of the switches is that they provide electrical isolation across floors and bandwidth can be added to individual switches when necessary. This will be handy for streaming video usage, etc. Cisco will be sending the order info on the switches. Judy should have her estimate in on the project sometime this week.
  2. The E-Reserves machines & scanners order has gone out.
  3. Erich has received his replacement scanners.
  4. Erich is requesting new CD ROM burner software.
  5. The new projector recently received for the East conference room will be housed here in Systems and made available for check-out only in this building. It requires hookup with some sort of computer.
Networking
  1. IPX problems continue - room 206 MSL has the problem where the slower machines can get to Novell, but the faster machines are not able! We'll search around for printers using the Novell protocol and turn it off - to see if that makes any difference. The IPX problems in West & the rest of MSL have subsided.
  2. Proxy (Squid) server - a change was made so the logs will auto rotate. It will keep nine days worth of logs. We are still getting lots of outsiders trying to use the server to get to ad & some porn sites. The server has been blocked, but they continue to try.
  3. West wiring - they're finishing up the first floor, the whole job should be completed by mid-May.
  4. NT10 - the testing of it was not successful. The software had past its expiration date. The Novell server came back up, but the date reverted to 1988. Whenever it goes down, the date has to be reset by hand. NT10 is currently not operating, but Will is working on it.
  5. BackUp Exec - will be used to copy individual mailboxes.
  6. Journalism - waiting to hear from people over there regarding IP requests. Will need a patch cable for the fiber wiring. The AP wire service will require large screens and video cards. Journalism will have to deal with Xerox regarding their printing needs.
Software Projects Status
  1. Film Log Q&A - Winston & Hilton demoed it this morning. There were some permissions problems. This is version 2. They are happy with the way it looks, but are requesting some adjustments. They will meet again next Monday.
  2. Ying is working on a Historical Log for time spent on software projects. Will be designed similar to PAULA.
  3. Systems wiring database - Bobbie has a few changes she needs to get to Ying. She is also requesting that a print capability be added.
  4. Suzy has been investigating the Exchange Help system and was able to set up the Resource Services conference room & laptop to be set up on the calendar, as requested by Diana Hagan. It worked so well that we may set up the Systems laptops for that type of checkout. Procedures would be required. Suzy will get together with Debra regarding this matter.
  5. While lurking around in the Outlook applications, Suzy found that it has a Trouble Ticket demo in it. She thinks it could work in our environment. It auto notifies and forwards information - complete with details. Suzy will get together with Debra on this, too.
DP Coordinator
  1. Wednesday is the last day of classes for the students. We won't have any more students until the beginning of summer semester - about 3 weeks.
  2. Lana's last day was this past Thursday.
  3. Melissa will be gone for the summer & a new student is in the process of being hired.
  4. Lindsay will be here until she leaves for England at the beginning of June. We'll need to hire a new student for her position.
Calendar
  1. Rich is out all this week, April 24-28.
  2. Bill will be out of the office this afternoon.
  3. Bobbie will be out April 27-28.
  4. Winston will be out June 19-23.
  5. Bobbie will be out July 10-21.
  6. Will is scheduled to be out July 24-28.


Back to the Systems Homepage

Last updated May 1, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu